If the interest rate is 4 percent, what is the present value of a bond that matures in three years, pays

$80 one year from now, $80 two years from now, and $1,080 three years from now?
Answer #1

ANSWER:

I = 4%

PV = cash flow in year 1(p/f,i,n) + cash flow in year 2(p/f,i,n) + cash flow in year 3(p/f,i,n)

PV = 80(p/f,4%,1) + 80(p/f,4%,2) + 1,080(p/f,4%,3)

PV = 80 * 0.9615 + 80 * 0.9246 + 1,080 * 0.889

PV = 76.92 + 73.968 + 960.12

PV = 1,111.008

so the present value of the bond is $1,111